### v1

# arXiv:2603.01359v1[eess.SY] 02 Mar 2026

Abstract
Over the last decades, progress in modal analysis has enabled increasingly routine use of modal parameters, including those extracted from in-situ measurements, for applications such as structural health monitoring and finite element model updating. For output-only identification, or Operational Modal Analysis (OMA), widely adopted approaches include Stochastic Subspace Identification (SSI) methods and the Natural Excitation Technique combined with the Eigensystem Realization Algorithm (NExT-ERA). Nevertheless, SSI-based techniques may become cumbersome on large systems, while NExT-ERA fitting can struggle when measurements are contaminated by noise. To alleviate these, this work investigates an OMA frequency-domain formulation for aeronautical structures by coupling the Loewner Framework (LF) with NExT, yielding the proposed NExT-LF method. The method exploits the computational efficiency of LF together with the impulse response function retrieval enabled by NExT. NExT-LF is assessed on two experimental benchmarks: the eXperimental BeaRDS 2 high-aspect-ratio wing main spar and an Airbus Helicopters H135 bearingless main rotor blade. The identified modal parameters are compared against available experimental references and results obtained via SSI with Canonical Variate Analysis and NExT-ERA. The results show that the modes identified by NExT-LF correlate well with benchmark data, particularly for high-amplitude tests and in the low-frequency range.
keywords:
Tangential Interpolation; Operational Modal Analysis; Loewner Framework, High-Aspect-Ratio Wing; Aeronautical Structure; Natural Excitation Technique; Wing Spar; Helicopter Blade; Airbus Helicopters H135.
[…]
1 Introduction
Modal analysis is a key tool for characterising the structural dynamics of engineering systems Ewins (2000). It can be performed in an input-controlled experimental setting or by measuring a structure in situ, either during operation or at rest (Ambient Vibration Testing – AVT), without the application of an external controlled input. The former is referred to as Experimental Modal Analysis (EMA), whereas the latter is known as Operational Modal Analysis (OMA) Brincker and Ventura (2015). In OMA, which is the focus of this work, the objective is to extract modal parameters Reynders (2012)—natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$)—to describe the dynamic behaviour of a structure for purposes including Finite Element Model (FEM) updating Zanotti Fragonara et al. (2017) and Structural Health Monitoring (SHM) Cadoret et al. (2025). In aeronautical and aerospace engineering, OMA is employed in a broad range of applications. For instance, Jelicic et al. (2014) uses OMA results to track flutter stability in real time for a wind-tunnel wing model, while Sinske et al. (2017) reports in-flight aeroelastic identification for the German Aerospace Center (DLR) HALO research aircraft. OMA has also been applied during launch-vehicle rollout operations, as reported for Artemis I in Akers et al. (2025), and to extract modal characteristics from in-flight data for launch vehicles and missiles in Eugeni et al. (2018) and James et al. (1994), respectively. Rotorcraft applications are also well established: Ameri et al. (2013) applies OMA to Ground Vibration Testing (GVT) data from a Westland Lynx Mark 7 helicopter airframe, while Sibille et al. (2023) and Agneni et al. (2010) identify time-expired helicopter blades from an Airbus Helicopters H135 and an Agusta-Bell AB 204, respectively.
Traditional OMA techniques include the Natural Excitation Technique (NExT) paired with the Eigensystem Realization Algorithm (ERA), commonly denoted as NExT-ERA James III et al. (1993), Stochastic Subspace Identification with Canonical Variate Analysis (SSI-CVA) Van Overschee and De Moor (1996), and Frequency Domain Decomposition (FDD) Brincker et al. (2001). NExT-ERA and SSI-CVA are time-domain approaches, whereas FDD operates directly in the frequency domain. Recent developments have aimed at improving robustness to measurement noise and supporting consistent mode selection. For example, O’Connell and Rogers (2024); O’Connell et al. (2025) propose a robust COVariance-driven SSI (SSI-COV) formulation incorporating a probabilistic criterion to reduce the likelihood that outliers are misinterpreted as physical modes. With a related objective, Amador and Brincker (2024) introduces a subspace-based poly-reference complex frequency approach to reduce uncertainties associated with the null space of the system matrices used in the identification process, by adopting a subspace-driven parametric formulation and estimating modal properties via eigensystem realisations across selected subspaces. Another SSI-based development couples SSI-COV with Fuzzy C-Means (SSI-FCM) clustering to reject spurious modes in stabilisation diagrams, and has been demonstrated on a train bogie and a five-storey experimental structure in Lv et al. (2026). Furthermore, the Multi-Reference Transmissibility Complex Frequency method is proposed in Tarinejad and Amanzad (2026) to address bias errors related to the application of the Fast Fourier Transform (FFT) to non-stationary signals, although validation is limited to numerical systems. Finally, NExT combined with Dynamic Mode Decomposition (DMD) is introduced for SHM of aeronautical structures in Wu et al. (2026), building on the DMD formulation for experimental modal analysis in Saito and Kuno (2020), which in turn extends the original DMD framework for extracting coherent features in fluid flows Schimid (2010).
In parallel, automated OMA frameworks (AOMA) have been proposed to streamline the identification process, including machine learning-based strategies Mugnaini et al. (2022), hierarchical clustering Neu et al. (2017), k-means clustering de Almeida Cardoso et al. (2018), and Density-Based Spatial Clustering of Applications with Noise (DBSCAN) Sibille et al. (2023). AOMA is outside the scope of this work, and the interested reader is referred to Mostafaei and Ghamami (2025) for a recent review.
Within this broader context, the frequency-domain Loewner Framework (LF) has been considered for EMA Dessena et al. (2023) and subsequently extended to output-only identification in civil engineering by pairing it with NExT, yielding the NExT-LF approach Dessena et al. (2025). The aim of this coupling is to exploit the computational efficiency of the LF fitting processs while retaining the output-only capability provided by NExT. These characteristics make NExT-LF attractive for aeronautical applications, where datasets can be large and measurements may be affected by operational noise. To address this gap, the present work investigates NExT-LF on two aeronautically relevant experimental datasets: the eXperimental BeaRDS 2 (XB-2) flexible wing spar model Dessena et al. (2022) and an Airbus Helicopters H135 Bearingless Main Rotor (BMR) blade Weber et al. (2021). The two datasets are selected due to their different levels of complexity in terms of system dimensions (1.45 m vs. 4.99 m span) and test conditions (shaker excitation vs. AVT). The identification accuracy of NExT-LF is evaluated against concurrent identifications obtained via SSI-CVA and NExT-ERA, as well as against available benchmark results. The overall objective is to assess whether the accuracy achieved by NExT-LF is suitable for aeronautical applications. To achieve this aim, the remainder of the paper is structured as follows:
  • Section˜2 (Methodology) introduces the LF identification procedure, the output-only enabling method (NExT), and their coupling into NExT-LF;
  • Section˜3 (The eXperimental BeaRDS 2 Flexible Wing Spar Model) presents the XB-2 dataset and discusses the corresponding identification results;
  • Section˜4 (Airbus Helicopters H135 Bearingless Main Rotor Blade) reports the blade dataset and the associated modal identification outcomes;
  • Section˜5 (Discussion) discusses the main findings, limitations, and practical implications of applying NExT-LF to aeronautical structures;
  • Section˜6 (Conclusions) summarises the main conclusions and closes this article.
2 Methodology
2.1 Loewner Framework
The LF algorithm has previously been adopted for the modelling of multi-port electrical systems Lefteriu and Antoulas (2009) and, in the field of aeroservoelastic modelling, for aerodynamic model order reduction Quero et al. (2019). Subsequently, the first and second authors employed the LF to identify modal parameters in SIMO mechanical systems Dessena et al. (2023), demonstrated its computational efficiency Dessena et al. (2023), and examined its robustness to measurement noise for SHM applications Dessena et al. (2024). More recent work has extended the LF to enable the extraction of modal parameters from multi-input multi-output Dessena and Civera (2025); Dessena et al. (2026) and output-only systems Dessena et al. (2025). The present study considers this latter output-only formulation.
To introduce the LF properly, we first define the Loewner matrix $\bm{\mathbb{L}}$}: Let ($\mu_{j}$${v}_{j}$), j=1j=1,…,$q$ denote a row-wise collection of complex pairs and ($\lambda_{i}$${w}_{j}$), i=1i=1,…,$k$ a column-wise collection of complex pairs, with all points $\lambda_{i}$ and $\mu_{j}$ mutually distinct. The matrix $\bm{\bm{\mathbb{L}}}$, commonly termed the divided-differences $\bm{\mathbb{L}}$} matrix, is then given by:
$\bm{\bm{\mathbb{L}}}=\begin{bmatrix}\frac{\bm{v}_{1}-\bm{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\bm{v}_{1}-\bm{w}_{k}}{\mu_{1}-\lambda_{k}}\\ \vdots&\ddots&\vdots\\ \frac{\bm{v}_{q}-\bm{w}_{1}}{\mu_{q}-\lambda_{1}}&\cdots&\frac{\bm{v}_{q}-\bm{w}_{k}}{\mu_{q}-\lambda_{k}}\\ \end{bmatrix}\>\in\mathbb{C}^{q\times k}$ (1)
If there is a known underlying function $\boldsymbol{\phi}$, then $\bm{w}_{i}=\boldsymbol{\phi}(\lambda_{i})$ and $\bm{v}_{j}=\boldsymbol{\phi}(\mu_{j}).$
Karl Löwner first linked $\bm{\mathbb{L}}$ to rational interpolation, a setting that is frequently described as Cauchy interpolation Löwner (1934). Through this link, interpolants may be characterised via the determinants of suitable submatrices of $\bm{\mathbb{L}}$. Moreover, as shown in Antoulas et al. (2017); Mayo and Antoulas (2007), one can obtain rational interpolants directly from $\bm{\mathbb{L}}$. In the present work, the formulation adopted is the Loewner pencil, defined by the pair of matrices $\bm{\mathbb{L}}$ and $\bm{\mathbb{L}}_{s}$. The latter, $\bm{\mathbb{L}}_{s}$, is the shifted Loewner matrix, introduced later on.
To illustrate the working principle of the LF, consider a linear time-invariant dynamical system $\bm{\Sigma}$ characterised by $m$ inputs, $p$ outputs, and $k$ internal variables, expressed in descriptor form as:
$\bm{\Sigma}:\;\bm{E}\frac{d}{dt}\bm{x}(t)=\bm{A}\bm{x}(t)+\bm{B}\bm{u}(t);\;\;\;\bm{y}(t)=\bm{C}\bm{x}(t)+\bm{D}\bm{u}(t)$ (2)
where $\bm{x}(t)\in\mathbb{R}^{k}$ denotes the vector of internal variables, $\bm{u}(t)\in\mathbb{R}^{m}$ is the input signal, and $\bm{y}(t)\in\mathbb{R}^{p}$ is the corresponding output vector. The system is specified by the following constant matrices:
$\bm{E},\bm{A}\in\mathbb{R}^{k\times k},\;\bm{B}\in\mathbb{R}^{k\times m}\;\bm{C}\in\mathbb{R}^{p\times k}\;\bm{D}\in\mathbb{R}^{p\times m}$ (3)
The Laplace-domain transfer function $\bm{H}(s)$ corresponding to $\bm{\Sigma}$ may be expressed as a $p\times m$ rational matrix function, provided that the pencil $\bm{A}-\lambda\bm{E}$ is regular, i.e., $\bm{A}-\lambda\bm{E}$ is non-singular for at least one finite $\lambda\in\mathbb{C}$:
$\bm{H}(s)=\bm{C}(s\bm{E}-\bm{A})^{-1}\bm{B}+\bm{D}$ (4)
Let us consider the general setting of tangential interpolation, which is commonly described as rational interpolation along tangential directions Kramer and Gugercin (2016). The associated right interpolation dataset is defined as:
$\begin{gathered}(\lambda_{i};\bm{r}_{i},\bm{w}_{i}),\>i=1,\dots,\rho\\ \begin{matrix}\bm{\Lambda}=\text{diag}[\lambda_{1},\dotsc,\lambda_{k}]\in\mathbb{C}^{\rho\times\rho}\\ \bm{R}=[\bm{r}_{1}\;\dotsc\bm{r}_{k}]\in\mathbb{C}^{m\times\rho}\\ \bm{W}=[\bm{w}_{1}\;\dotsc\;\bm{w}_{k}]\in\mathbb{C}^{p\times\rho}\end{matrix}\Bigg\}\end{gathered}$ (5)
Similarly, the left interpolation data can be outlined:
$\begin{gathered}(\mu_{j},\bm{l}_{j},\bm{v}_{j}),\>j=1,\dots,v\\ \begin{matrix}\bm{M}=\mathrm{diag}[\mu_{1},\dotsc,\mu_{q}]\in\mathbb{C}^{v\times v}\\ \bm{L}^{T}=[\bm{l}_{1}\;\dotsc\bm{l}_{v}]\in\mathbb{C}^{p\times v}\\ \bm{V}^{T}=[\bm{v}_{1}\;\dotsc\;\bm{v}_{q}]\in\mathbb{C}^{m\times v}\end{matrix}\Bigg\}\end{gathered}$ (6)
The values $\lambda_{i}$ and $\mu_{j}$ denote the sampling points at which $\bm{H}(s)$ is queried, which in the present context coincide with the chosen frequency bins. The vectors $\bm{r}_{i}$ and $\bm{l}_{j}$ specify the right and left tangential directions and are often drawn at random in practical implementations Quero et al. (2019). The quantities $\bm{w}_{i}$ and $\bm{v}_{j}$ are the corresponding tangential measurements. The rational interpolation task is then satisfied by enforcing consistency between $\bm{w}_{i}$, $\bm{v}_{j}$ and the transfer function $\bm{H}$ associated with the realisation $\bm{\Sigma}$ in Equation˜2:
$\bm{H}(\lambda_{i})\bm{r}_{i}=\bm{w}_{i},\>j=1.\dots,\rho\;\;\text{and}\;\;\bm{l}_{i}\bm{H}(\mu_{j})=\bm{v}_{j},\>i=1,\dots,v$ (7)
so that the Loewner pencil complies with Equation˜7. Next, consider a collection of points $Z={z_{1},\ldots,z_{N}}\subset\mathbb{C}$ together with a rational function $\bm{y}(s)$. For each $i=1,\ldots,N$, define the sampled values $\bm{y}_{i}:=\bm{y}(z_{i})$, and gather them in the set $Y={\bm{y}_{1},\ldots,\bm{y}_{N}}$. Upon introducing the corresponding left and right partitions of the data, one obtains the following relations:
$\begin{split}Z=\{\lambda_{1},\dots,\lambda_{\rho}\}\cup\{\mu_{1},\dots,\mu_{v}\}\;\;\text{and}\;\;Y=\{\bm{w}_{1},\dots,\bm{w}_{\rho}\}\cup\{\bm{v}_{1},\dots,\bm{v}_{v}\}\end{split}$ (8)
where $N=p+v$. Consequently, the Loewner matrix $\bm{\mathbb{L}}$ can be written in the form:
$\bm{\bm{\mathbb{L}}}=\begin{bmatrix}\frac{\bm{v}_{1}\bm{r}_{1}-\bm{l}_{1}\bm{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\bm{v}_{1}\bm{r}\rho-\bm{l}_{1}\bm{w}\rho}{\mu_{1}-\lambda\rho}\\ \vdots&\ddots&\vdots\\ \frac{\bm{v}_{v}\bm{r}_{1}-\bm{l}_{v}\bm{w}_{1}}{\mu_{v}-\lambda_{1}}&\cdots&\frac{\bm{v}_{v}\bm{r}\rho-\bm{l}_{v}\bm{w}\rho}{\mu_{v}-\lambda\rho}\\ \end{bmatrix}\>\in\mathbb{C}^{v\times\rho}$ (9)
Noting that $\bm{v}_{v}\bm{r}_{p}$ and $\bm{l}_{v}\bm{w}_{p}$ are scalar quantities, the Sylvester relation associated with $\bm{\mathbb{L}}$ can be written as:
$\bm{\bm{\mathbb{L}}}\bm{\Lambda}-\bm{M}\bm{\bm{\mathbb{L}}}=\bm{L}\bm{W}-\bm{V}\bm{R}$ (10)
The shifted Loewner matrix $\bm{\mathbb{L}}_{s}$ is obtained by constructing the Loewner matrix for the function $s\bm{H}(s)$, namely:
$\bm{\bm{\mathbb{L}}}_{s}=\begin{bmatrix}\frac{\mu_{1}\bm{v}_{1}\bm{r}_{1}-\lambda_{1}\bm{l}_{1}\bm{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\mu_{1}\bm{v}_{1}\bm{r}_{\rho}-\lambda_{\rho}\bm{l}_{1}\bm{w}_{\rho}}{\mu_{1}-\lambda_{\rho}}\\ \vdots&\ddots&\vdots\\ \frac{\mu_{v}\bm{v}_{v}\bm{r}_{1}-\lambda_{1}\bm{l}_{v}\bm{w}_{1}}{\mu_{v}-\lambda_{1}}&\cdots&\frac{\bm{v}_{v}\bm{r}_{\rho}-\bm{l}_{v}\bm{w}_{\rho}}{\mu_{v}-\lambda_{\rho}}\\ \end{bmatrix}\>\in\mathbb{C}^{v\times\rho}$ (11)
In an analogous manner, the corresponding Sylvester relation takes the form:
$\bm{\bm{\mathbb{L}}}_{s}\Lambda-\bm{M}\bm{\bm{\mathbb{L}}}_{s}=\bm{L}\bm{W}\bm{\Lambda}-\bm{M}\bm{V}\bm{R}$ (12)
Without loss of generality, $\bm{D}$ may be set to zero, since within the LF framework its contribution does not influence the tangential interpolation conditions Mayo and Antoulas (2007). Accordingly, Equation˜4 simplifies to:
$\bm{H}(s)=\bm{C}(s\bm{E}-\bm{A})^{-1}\bm{B}$ (13)
A realisation of minimal order is attainable only when the underlying system is both controllable and observable. Under the assumption that the available samples originate from a system whose transfer function is given by Equation˜13, the generalised tangential observability matrix $\mathcal{O}_{v}$ and the generalised tangential controllability matrix $\mathcal{R}_{\rho}$ are introduced in Lefteriu and Antoulas (2010). It then follows that Equation˜9 and Equation˜11 may be recast in the form:
$\bm{\bm{\mathbb{L}}}=-\mathcal{O}_{v}\bm{E}\mathcal{R}_{\rho}$ $\bm{\bm{\mathbb{L}}}_{s}=-\mathcal{O}_{v}\bm{A}\mathcal{R}_{\rho}$ (14)
Next, define the Loewner pencil as regular by requiring that its generalised eigenvalues do not coincide with the sampling points, i.e., $\mathrm{eig}((\bm{\bm{\mathbb{L}}},\bm{\bm{\mathbb{L}}}_{s}))\neq(\mu_{i},\lambda_{i})$:
$\bm{E}=-\bm{\bm{\mathbb{L}}},$ $\bm{A}=-\bm{\bm{\mathbb{L}}}_{s},$ $\bm{B}=\bm{V},$ $\bm{C}=\bm{W}$ (15)
Accordingly, the resulting rational interpolant may be written as:
$\bm{H}(s)=\bm{W}(\bm{\bm{\mathbb{L}}}_{s}-s\bm{\bm{\mathbb{L}}})^{-1}\bm{V}$ (16)
The preceding derivation is restricted to the minimal-data setting, a situation that is rarely met in practice. However, the LF framework admits a natural extension that can incorporate redundant sampling points in an efficient manner. To that end, we introduce the following assumption:
$\begin{split}\mathrm{rank}[\zeta\bm{\bm{\mathbb{L}}}-\bm{\bm{\mathbb{L}}}_{s}]=\mathrm{rank}[\bm{\bm{\mathbb{L}}}\>\bm{\bm{\mathbb{L}}}_{s}]=\mathrm{rank}\begin{bmatrix}\bm{\bm{\mathbb{L}}}\bm{\bm{\mathbb{L}}}_{s}\end{bmatrix}=k,\;\forall\zeta\in\{\lambda_{j}\}\cup\{\mu_{i}\}\end{split}$ (17)
Subsequently, a truncated Singular Value Decomposition (SVD) is computed for the matrix $\zeta\bm{\mathbb{L}}-\bm{\mathbb{L}}_{s}$:
$\textrm{svd}(\zeta\bm{\bm{\mathbb{L}}}-\bm{\bm{\mathbb{L}}}_{s})=\bm{Y}\bm{\Sigma}_{l}\bm{X}$ (18)
where $\mathrm{rank}(\zeta\bm{\bm{\mathbb{L}}}-\bm{\bm{\mathbb{L}}}_{s})=\mathrm{rank}(\bm{\Sigma}_{l})=\mathrm{size}(\bm{\Sigma}_{l})=k,\bm{Y}\in\mathbb{C}^{v\times k}$ and $\bm{X}\in\mathbb{C}^{k\times\rho}$. Note that:
$\begin{split}-\bm{A}\bm{X}+\bm{E}\bm{X}\bm{\Sigma}_{l}=\bm{Y}^{*}\bm{\bm{\mathbb{L}}}_{s}\bm{X}^{*}\bm{X}-\bm{Y}^{*}\bm{\bm{\mathbb{L}}}\bm{X}^{*}\bm{X}\bm{\Sigma}_{l}=\bm{Y}^{*}(\bm{\bm{\mathbb{L}}}_{s}-\bm{\bm{\mathbb{L}}}\bm{\Sigma}_{l})=\bm{Y}^{*}\bm{V}\bm{R}=\bm{BR}\end{split}$ (19)
In an analogous manner, the identity $\bm{Y}\bm{A}+\bm{M}\bm{Y}\bm{E}=\bm{L}\bm{C}$ is obtained, where $\bm{X}$ and $\bm{Y}$ act as the generalised controllability and observability matrices of the system $\bm{\Sigma}$, respectively, under the assumption $\bm{D}=0$.
Having verified that both the right and left interpolation conditions are fulfilled, the Loewner realisation accommodating redundant data may be expressed as:
$\begin{split}\bm{E}=-\bm{Y}^{*}\bm{\bm{\mathbb{L}}}\bm{X},\quad\quad\bm{A}=-\bm{Y}^{*}\bm{\bm{\mathbb{L}}}_{s}\bm{X,}\quad\quad\bm{B}=\bm{Y}^{*}\bm{V},\quad\quad\bm{C}=\bm{W}\bm{X}\end{split}$ (20)
The expression in Equation˜20, which specifies the Loewner realisation in the presence of redundant data, is adopted as the reference formulation in the remainder of this work. The complete step-by-step account procedure is provided in Mayo and Antoulas (2007); Antoulas et al. (2017). Finally, the modal parameters of the system may be obtained by performing an eigenanalysis of the matrices $\bm{A}$ and $\bm{C}$ appearing in Equation˜20. In practical modal analysis, this procedure is repeated over a range of model orders $k$, selected to capture the minimum set of modes of interest while identifying the highest order for which these modes remain stably estimated. From a practical standpoint, the frequency domain data, usually an FRF, $\bm{H}_{real}(s)$ is fed to the LF, which then breaks down the data into two subsets, the left and right data, so that the system matrices of the transfer function in Equation˜13 can be retrieved.
2.2 Natural Excitation Technique
NExT enables modal identification from output-only measurements by exploiting the statistical structure of ambient (unmeasured) excitations. In many operational settings, the applied forces cannot be recorded reliably (e.g. wind, traffic, or in-service aerodynamic loads), which prevents the direct construction of FRFs from measured inputs and outputs. NExT circumvents this limitation by assuming the unknown excitation as a random process with sufficiently broad frequency content to excite the modes of interest, and by using output correlations as impulse-response-type quantities James III et al. (1993), the system IRFs can be obtained.
Under the standard assumptions of linearity and time invariance, the output-only problem can be interpreted through a stochastic state-space model in discrete time, an LTI:
$x_{k+1}=Ax_{k}+w_{k},\qquad y_{k}=Cx_{k}+v_{k},$ (21)
where $k\in\mathbb{N}$ is the sample index, $x_{k}\in\mathbb{R}^{n}$ is the state vector, $y_{k}\in\mathbb{R}^{p}$ is the measured output vector, and $A\in\mathbb{R}^{n\times n}$ and $C\in\mathbb{R}^{p\times n}$ are the state-transition and output matrices, respectively. The terms $w_{k}$ and $v_{k}$ represent process and measurement noise. NExT typically assumes that the effective excitation is zero-mean and stationary over the analysed record, and that it can be treated as broadband (often idealised as white noise) within the frequency range of interest. Marked departures from stationarity (e.g. strongly time-varying operating conditions) reduce the reliability of correlation estimates and may bias the inferred impulse responses.
Let $y_{i}(t)$ and $y_{j}(t)$ denote two measured responses (e.g. accelerations) acquired at distinct sensor locations. The cross-correlation function is defined as
$R_{y_{i}y_{j}}(\tau)=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}y_{i}(t)\,y_{j}(t+\tau)\,\mathrm{d}t,$ (22)
where $\tau$ is the time lag. For finite records, (22) is estimated from sampled data. Using $N$ samples $y_{i}[k]$ and $y_{j}[k]$, one common unbiased estimator for non-negative lags is
$\widehat{R}_{y_{i}y_{j}}[\ell]=\frac{1}{N-\ell}\sum_{k=0}^{N-\ell-1}y_{i}[k]\,y_{j}[k+\ell],\qquad\ell=0,1,\dots,N-1.$ (23)
When the ambient excitation is broadband and the system is LTI, the correlation functions can be written as a superposition of decaying sinusoids whose damped frequencies and decay rates coincide with those of the structural modes James III et al. (1993). In practice, this means that $\widehat{R}_{y_{i}y_{j}}[\ell]$ (with $\ell$ denoting the discrete lag index, i.e. $\tau=\ell\Delta t$ for sampling period $\Delta t$) has the same functional form as a free-decay (IRF) for modal identification purposes, up to an unknown scaling and sensor-pair-dependent phase. A convenient representation is:
$\widehat{R}_{y_{i}y_{j}}(\tau)\approx\sum_{r=1}^{n_{m}}A_{r,ij}\,\mathrm{e}^{-\zeta_{r}\omega_{r}\tau}\sin\!\bigl(\omega_{d,r}\tau+\theta_{r,ij}\bigr),\qquad\omega_{d,r}=\omega_{r}\sqrt{1-\zeta_{r}^{2}},$ (24)
where $n_{m}$ is the number of participating modes, $\omega_{r}$ and $\zeta_{r}$ are the natural frequency and damping ratio of mode $r$, and $A_{r,ij}$ and $\theta_{r,ij}$ are constants that depend on the sensor pair and on the (unknown) spatial distribution of the excitation.
Given $p$ output channels, NExT forms a set of impulse-response-like sequences by selecting a reference channel $j$ and computing $\widehat{R}_{y_{i}y_{j}}[\ell]$ for $i=1,\dots,p$. The resulting collection $\{\widehat{R}_{y_{1}y_{j}},\widehat{R}_{y_{2}y_{j}},\dots,\widehat{R}_{y_{p}y_{j}}\}$ is interpreted as a multiple-output IRF associated with the chosen reference. An equivalent frequency-domain route, not considered in this work, computes cross-spectral densities and then obtains correlations via an inverse Fourier transform. Both approaches are consistent provided the stationarity assumptions are satisfied and the same normalisation conventions are used.
The performance of NExT depends on (i) approximate stationarity of the operating condition over the analysed record, (ii) sufficient bandwidth of the effective excitation to activate the modes of interest, and (iii) adequate signal-to-noise ratio across the sensor set. Close modes, strong tonal components (harmonics), or dominant narrow-band excitations can complicate the separation of modal contributions and may require longer records, additional averaging, or complementary processing (e.g. harmonic removal).
2.3 NExT-LF
Summarising, NExT can retrieve IRFs from output-only data. IRFs are the time domain counterpart of the FRFs, such that they can be brought into the frequency domain by means of an FTT and then become a suitable input for frequency domain methods, such as the LF. In this work, NExT-LF follows the workflow shown in Figure˜1.
Figure 1: NExT-LF workflow.
3 The eXperimental BeaRDS 2 Flexible Wing Spar Model
[…]
Figure 2: XB-2 wing main spar geometrical characteristics and corresponding accelerometer locations ( ). The odd-numbered accelerometers are positioned on the positive $z$ side, whereas the even-numbered sensors are mounted on the negative $z$ side (retrieved from Dessena et al. (2025a)).
The experimental dataset employed in this work is sourced from Dessena et al. (2022) [^1] The experimental and benchmark data can be retrieved from https://doi.org/10.17862/cranfield.rd.19077023, which also reports benchmark EMA results. The spar is excited for 20 m using a bandwidth-limited random input in the range 2-400 Hz, with an RMS amplitude of 0.305 g. The excitation is applied using a Data Physics Signal Force™ modal shaker operated under closed-loop control via DP760™ software. As shown in Figure˜2, the accelerometers are arranged in a 4 by 2 grid to measure vertical acceleration, thereby capturing both flapwise (out-of-plane bending in the $x$-$z$ plane) and pitching (torsion around the $x$ axis) responses. The signals were sampled at $f_{s}=$ 5120 Hz using a National Instruments cDAQ-9178 and saved via an in-house LabVIEW programme [^2] At the time the first author was affiliated with Cranfield University.. The recorded acceleration time series are then converted to ms-2, band-pass filtered between 3.25 and 85 Hz to suppress low-frequency drift, and band-stop filtered between 49.5 and 50.5 Hz to remove mains-related interference [^3] In the UK, the mains electricity AC frequency is 50 Hz.. The upper passband limit is set to 85 Hz, since all modes of interest lie below this frequency. Further details on the experimental setup and instrumentation used in the test are avaible in Dessena et al. (2022) and additional details on a nonlinear vibration testing campaign on XB-2 and components can be found in Dessena et al. (2025b).
Using the eight output-channel time histories, the response spectra can be estimated in terms of power spectral density (PSD). This is computed in MATLAB via the built-in pwelch function [^4] https://uk.mathworks.com/help/signal/ref/pwelch.html. Prior the PSD estimation, the signal is downsampled to 256 Hz. Then, the resulting PSD estimates, as shown in Figure˜3, are obtained by adopting 2048 discrete Fourier transform points, an overlap of 256 samples, and a Hamming window of length 1024 to reduce variance and provide a smoother spectral estimate.
Figure 3: PSDs of the eight acceleration response channels measured on the XB-2 wing main spar (retrieved from Dessena et al. (2025a)).
As indicated by Figure˜3, the PSD exhibits three dominant peaks, which can be attributed to the three modes of interest known to occur at approximately 4.8, 27, and 76.8 Hz. The pronounced notch near 50 Hz arises from the applied band-stop filter. In addition, within the 60-70 Hz range, a resonance-like feature and an anti-resonance-like shape (most evident in channels #3 and #4, respectively) are observed. These are associated with local interactions involving the spar reinforcement plates to which the accelerometers are mounted, rather than with the global structural dynamics. Consistently, these features are absent from the PSDs of the remaining channels and do not appear in any FRF. The FRFs are omitted here for brevity, but are reported in the benchmark study Dessena et al. (2022).
After this step, the identification procedure begins by providing NExT with the eight measured output acceleration time series, from which IRFs discretised over 2048 points are obtained. In this work, the acceleration signal from accelerometer #8 (see Figure˜2) is used as the sole reference channel for the NExT algorithm. This choice is motivated by the observation that the corresponding NExT-derived FRFs, as shown in Figure˜4), present the three expected resonance peaks more clearly (higher amplitude) than when other reference channels are used.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
Figure 4: Equivalent FRFs obtained from the NExT-derived IRF of the XB-2 wing spar. (a) to (h) show the FRFs obtained, respectively, by setting output channel signals 1 to 8 the reference signal for NExT.
These FRFs are then used as input to the LF for modal parameter estimation. In the NExT-ERA workflow, the frequency-domain step is not required, since the IRFs are provided directly to the ERA algorithm to extract the modal parameters. For SSI, modal parameters are estimated by supplying the measured acceleration time series directly to the method.
To limit the influence of outliers and non-physical modes, the identification is repeated across multiple model orders $k$, with $k\in[6,50]$, and stabilisation diagrams are used to retain only stable modes. The range of $k$ was chosen to obtain a comparable count of stable modes across the methods. As discussed later, this condition could not be met for ERA. The stability assessment considers $\omega_{n}\in[0,83]\,\mathrm{Hz}$ with a stability parameter of $0.5\%$, and $\zeta_{n}\in[0.01,0.03]$ with a stability parameter of $5\%$. A mode is accepted as stable only when the criteria are satisfied for five consecutive identifications. $\bm{\phi}_{n}$ consistency is enforced by requiring a Modal Assurance Criterion (MAC) value not smaller than $0.95$ among the subsequent ($k$-wise) identified modes. The stabilisation diagrams for NExT-LF, NExT-ERA, and SSI-CVA are reported in Figure˜4.
((a))
((b))
((c))
Figure 5: Stabilisation diagrams obtained using NExT-LF (a), NExT-ERA (b), and SSI (c) for the modal identification of the XB-2 wing spar (retrieved from Dessena et al. (2025a)).
As shown in Figure˜5, the NExT-LF and SSI-CVA results exhibit markedly stronger stabilisation than the NExT-ERA case, with the difference being most evident in the damping estimates $\zeta_{n}$. In particular, NExT-ERA yields virtually no modes that stabilise in damping. This behaviour is consistent with the identified $\omega_{n}$ and $\zeta_{n}$ values reported in Tables˜1 and 2, respectively.
Table 1: Natural frequencies identified from the XB-2 main spar.
Natural Frequency [Hz] (difference w.r.t. to Benchmark [%])
Mode # Benchmark Dessena et al. (2022) NExT-LF NExT-ERA SSI
1 4.855 4.930 4.689 4.891
(-) (1.53) (-3.42) (0.73)
2 26.966 26.922 26.942 26.969
(-) (-0.16) (-0.09) (0.01)
3 76.851 77.011 76.923 76.940
(-) (0.21) (0.09) (0.12)
Table 2: Identified modal damping ratios for the XB-2 main spar.
Modal Damping Ratios [-] (difference w.r.t. to Benchmark [%])
Mode # Benchmark Dessena et al. (2022) NExT-LF NExT-ERA SSI
1 0.033 0.016 0.015 0.027
(-) (-51.17) (-53.00) (-17.01)
2 0.010 0.011 0.012 0.011
(-) (7.29) (11.40) (3.06)
3 0.014 0.015 0.013 0.014
(-) (7.35) (-7.46) (-3.17)
The $\omega_{n}$ identified using NExT-LF, NExT-ERA, and SSI-CVA are in close agreement with the benchmark results reported in Dessena et al. (2022), with the largest discrepancy being $3.42\%$ for $\omega_{1}$ obtained via NExT-ERA. In contrast, the $\zeta_{n}$ in Table˜2 show noticeably larger differences. Specifically, the absolute error in $\zeta_{1}$ exceeds $50\%$ for both NExT-LF and NExT-ERA, with NExT-LF providing a smaller deviation than NExT-ERA, whereas SSI-CVA limits the absolute error to approximately $17\%$. For $\zeta_{2}$ and $\zeta_{3}$, the NExT-LF estimates deviate by about $7.3\%$ in absolute terms, while NExT-ERA produces larger errors, namely $11.4\%$ and $7.46\%$, respectively. Overall, SSI-CVA yields damping estimates that are most consistent with the reference identification, although the absolute error for $\zeta_{1}$ remains above $17\%$.
After discussing the identifications of $\omega_{n}$ and $\zeta_{n}$, the results for the $\bm{\phi}_{n}$ are now considered. Consistently with the frequency estimates, the mode shapes identified from the output-only measurements show excellent agreement with the benchmark, with the diagonal MAC values being equal, or very close, to unity. Figure˜6 presents the $\bm{\phi}_{n}$ obtained via NExT-LF, NExT-ERA, and SSI from output-only data, overlaid with the reference $\bm{\phi}_{n}$ reported in Dessena et al. (2022). The identified $\bm{\phi}_{n}$ closely reproduce the benchmark results.
Figure 6: $\bm{\phi}_{n}$ of the XB-2 wing spae identified using the output-only methods, overlaid with the benchmark results (retrieved from Dessena et al. (2025a)).
4 Airbus Helicopters H135 Bearingless Main Rotor Blade
In order to generalise the findings of the previous sections, a second benchmark aeronautical structure should be considered. For this aim an Airbus Helicopters H135 BMR blade is chosen. The Airbus Helicopters H135, shown in Figure˜7(a), is a lightweight rotorcraft capable of a maximum forward flight speed of 136 kts, a maximum range of 342 nm, and a maximum take-off mass of 3100 kg. In this work, the focus is on its BMR blade, which is made of the pitch control cuff with an elastomeric lead-lag damper, an aerofoil section and a flexbeam. This arrangement allows it to incorporate the conventional blade mechanical flapping, lagging, and torsional hinges into a single structural arrangement. More information on the Airbus Helicopters H135 BMR blade is available in Bansemir and Müller (1997). From a geometric perspective, the BMR blade has an equivalent chord of 0.288 m and spans 4.99 m from the hub attachment. The H135 BMR blade specimen considered here, see Figure˜7(b) is an unserviceable [^5] In the aircraft maintenance sense. item, still in relatively good shape, which has been dynamically tested, via AVT, in the former Cranfield University project BladeSense Weber et al. (2021).
((a))
((b))
Figure 7: (a) Airbus Helicopters H135 helicopter in Western Counties Air Operations Unit livery (UK – retrieved from Felce (2011)) and (b) the test setup of the ambient vibration testing on the Airbus Helicopter H135 BMR blade (adapted from Weber et al. (2021)).
The AVT was carried out in the aeroelastic laboratory of the Sir Peter Gregson Aerospace Integration Research Centre at Cranfield University, under controlled temperature and humidity conditions. The stationary, i.e. not revolving, blade is clamped at its end and instrumented with 9 nine triaxial accelerometers (A1-9), recording in-plane ($z$-axis) and out-of-plane transverse ($y$-axis) acceleration response from ambient vibration666xx-axis accelerations are recorded from A9, but are disregarded in this study.. The accelerometer types are shown in Figure˜7(b) and their layout and spanwise position in Figure˜8. The accelerometers spanwise locations in Figure˜8 are determined graphically from the Figures in Weber et al. (2021). Additionally, note that   highlights the 4 PCB Model 356B18 accelerometers used, while   and   identify, respectively, the 3 PCB Model 356A16 and 2 PCB Model 356A45 accelerometers.
Figure 8: Sensors layout of the instrumented Airbus Helicopter H135 BMR blade: A9-6 PCB Model 356B18 ( ), A5-3 PCB Model 356A16 ( ), and A2-1 PCB Model 356A45 ( ). Note that the $z$-axis is positive in the out-of-page direction.
From the 9 accelerometers, 18 acceleration time series are recorded at a sampling frequency $f_{s}=$ 2560 Hz for 600 s, using a similar setup to what is described for the XB-2 spar in Section˜3, involving a National Instruments cDAQ-9178 and then in-house LabVIEW programme. Prior to identification, the signals are decimated to 256 Hz, demeaned, detrended, and bandpass-filtered between 0.4 and 100 Hz to exclude potential low- and high-frequency drifts with the bandpass [^7] https://uk.mathworks.com/help/signal/ref/bandpass.html built-in MATLAB function. From here, a PSD of the 18 channels and the ANPSD are found to qualitatively assess the quality of the response . The PSDs and ANPSD, shown in Figure˜9 are calculated using Welch’s method over 2048 discrete points with a Hamming window of length 2048 with an overlap of 256.
Figure 9: PSDs and ANPSD from the signals recorded from accelerometers A1-9 on the irbus Helicopter H135 BMR blade.
As it is clear from Figure˜9, this case poses a tougher identification challenge than the XB-2 spar, as the AVT amplitudes are much smaller than those produced by shaker excitation, resulting in lower modal amplitudes. As for the other cases, the identification procedure starts by obtaining the IRF from NExT, discretised over 4098 points, using two reference channels, A1 in $x$ and A4 in $x$. This is necessary in this case, as using a reference channel for each direction increases the likelihood of obtaining accurate modal parameter estimates Caicedo (2011). The obtained IRFs are fed directly into ERA, while they are transformed to the frequency domain via an FFT discretised over 2048 points for use with the LF. The SSI-based identification is carried out directly from the time-domain data.
The identification is carried at a given order range $k\in[22,130]$ and stabilisation diagrams are used to exclude spurious and non physical modes. For this aim, stabilisation parameters, remaining the same across NExT-LF, NExT-ERA, and SSI-CVA, are defined. The stabilisation assessment is carried out over $f_{n}\in[0,100]\,$ Hz and $\zeta_{n}\in[0.003,0.03]$. Frequency stability is enforced with a relative tolerance of $0.5\%$ (i.e., $\Delta\omega_{n}/\omega_{n}\leq 0.005$), while damping stability is checked with a $5\%$ tolerance (i.e., $\Delta\zeta/\zeta\leq 0.05$). An additional absolute criterion on the frequency variation is also applied, $|\Delta\omega_{n}|\leq\epsilon$, with $\epsilon=0.1$. A mode is accepted as stable only if these criteria are satisfied for three consecutive identifications. The consistency of the $\bm{\phi}_{n}$ is enforced by requiring a MAC not smaller than $0.95$ among the subsequent stable identified modes. The stabilisation diagrams obtained from the NExT-LF, NExT-ERA, and SSI-CVA identification reported above are shown in Figure˜10.
((a))
((b))
((c))
Figure 10: Stabilisation diagrams superimposed on the ANPSD and obtained using NExT–LF (a), NExT–ERA (b), and SSI (c) for the modal identification of the Airbus Helicopters H135 BMR blade.
The stabilisation diagrams in Figure˜10 show that NExT–ERA, as observed for the XB-2 spar, struggles to identify most of the system modes. By contrast, the stabilisation diagrams obtained with NExT–LF and SSI–CVA are clearer and more consistent, in terms of both stability and the modes identified. It is also worth noting that some modes appear to be identified only by one of the two methods, with each approach capturing modes that are not recovered by the other.
The extracted modal parameters, $\omega_{n}$ and $\zeta_{n}$, are presentend in Table˜3 with results from the literature. These include results from AOMA Mugnaini et al. (2022), containing also $\bm{\phi}_{n}$, and Weber et al. (2021), including $\omega_{n}$ from a calibrated FEM and $\omega_{n}$ and $\zeta_{n}$ from a separate experimental configuration, based on a series of GVTs conducted using a random-on-random excitation waveform. Please note that in Table˜3, flapping refers to out-of-plane bending (within the xxyy plane), lagging to in-plane bending (within the xxyy plane), pitching to torsional rotations about the $x$-axis, and coupled to a mode exhibiting a combination of two or more of the dominant motions described above.
Table 3: Natural frequencies and damping ratios obtained with different methods for the Airbus Helicopters H135 BMR blade. Values in parentheses indicate the relative difference (%) w.r.t. Mugnaini et al. (2022).
Mode # Natural Frequencies [Hz] – (Relative difference w.r.t. Mugnaini et al. (2022) [%]) Damping ratio [%] – (Relative difference w.r.t. Mugnaini et al. (2022) [%])
NExT-LF NExT-ERA SSI-CVA AOMA Mugnaini et al. (2022) Exp. Weber et al. (2021) FEM Weber et al. (2021) NExT-LF NExT-ERA SSI-CVA AOMA Mugnaini et al. (2022) Exp. Weber et al. (2021)
1 - 1st flapping 1.03 1.03 1.02 1.02 1.02 0.90 2.33 2.72 0.68 0.55 1.05
(0.57) (0.65) (0.10) (0.00) (-11.76) (324.22) (394.53) (23.07) (90.91)
2 - 1nd lagging 2.68 - - 2.71 2.60 3.41 2.76 - - 0.70 5.50
(-1.01) (-4.06) (25.83) (293.86) (685.71)
3 - 2nd flapping 5.44 5.44 5.44 5.43 5.39 4.98 0.55 0.60 0.76 0.50 1.13
(0.18) (0.13) (0.15) (-0.74) (-8.29) (10.43) (19.06) (51.41) (126.00)
4 - 3rd flapping 15.58 15.57 15.55 15.55 15.44 15.28 0.47 0.34 0.35 0.30 0.97
(0.18) (0.16) (0.03) (-0.71) (-1.74) (55.93) (13.85) (17.93) (223.33)
5 - 2nd lagging 17.83 17.90 17.93 18.11 18.68 22.72 1.91 1.49 1.48 1.12 -
(-1.55) (-1.18) (-1.01) (3.15) (25.46) (70.46) (33.42) (32.08)
6 - 1st coupled 19.69 18.90 18.98 - - - 0.87 2.60 2.62 - -
7 - 1st pitching 28.20 - 28.14 28.11 27.39 28.59 0.46 - 0.67 0.66 3.22
(0.31) (0.09) (-2.56) (1.71) (-31.03) (1.70) (387.88)
8 - 4th flapping 30.11 30.08 30.06 30.07 29.98 29.98 0.42 0.45 0.45 0.41 0.61
(0.15) (0.02) (-0.03) (-0.30) (6.65) (1.49) (10.90) (9.01) (48.78)
9 - 5th flapping 51.49 51.42 51.46 51.43 51.71 52.17 0.31 0.31 0.57 0.45 0.85
(0.12) (-0.03) (0.05) (0.54) (73.50) (-32.10) (-31.94) (26.72) (88.89)
10 - 3rd lagging 55.65 57.78 56.65 56.24 - - 2.64 1.58 2.68 2.52 -
(-1.05) (2.74) (0.74) (4.89) (-37.35) (6.20)
11 - 2nd coupled 59.14 60.27 59.50 - - - 1.08 1.13 2.34 - -
12 - 3rd coupled - - 68.09 - - - - - 1.65 - -
13 - 6th flapping 75.65 - 75.73 75.80 75.68 75.12 0.31 - 0.85 0.88 0.94
(-0.20) (-0.09) (-0.16) (-0.90) (-65.30) (-3.11) (6.82)
14 - 2nd pitching 82.05 - 81.86 81.71 81.63 81.80 1.04 - 1.60 1.82 1.50
(0.42) (0.19) (-0.10) (0.11) (-42.63) (-12.16) (-17.58)
From Table˜3, it is important to note that NExT-LF identifies two additional modes within the frequency range of interest, compared with the previous identification reported for AOMA Mugnaini et al. (2022) and the experimental results in Weber et al. (2021). These modes are located at $19.69~\mathrm{Hz}$ and $59.14~\mathrm{Hz}$, respectively. In addition, SSI-CVA identifies a further mode at $68.09~\mathrm{Hz}$. As shown later, these modes are all coupled in the $z$- and $y$-axis displacements, with the suspicion that the 2nd coupled mode may also include a torsional component; however, this cannot be fully confirmed due to the adopted instrumentation layout. Nevertheless, the coupled character of these modes provides a plausible explanation for why they are not identified by the simplified FEM reported in Weber et al. (2021). Furthermore, it is worth noting that NExT-LF finds more modes than NExT-ERA (13 vs 9), and the same number as SSI-CVA.
In terms of the identified natural frequencies $\omega_{n}$, the agreement across the different output-only identification methods is generally very good. With respect to the reference values obtained via AOMA Mugnaini et al. (2022), the relative differences remain below approximately $3\%$ for all coincident modes, the largest deviation being observed for NExT-ERA at the 10th mode. A similar level of agreement is observed when comparing with the experimental results reported in Weber et al. (2021). The consistency of the identified $\omega_{n}$ across NExT-LF, NExT-ERA, and SSI-CVA confirms the robustness of the proposed framework in terms of frequency estimation, even for higher-order flapping and lagging modes.
However, the same level of consistency is not observed for the damping ratios $\zeta_{n}$. In contrast to $\omega_{n}$, the identified $\zeta_{n}$ values exhibit a significantly larger dispersion across methods. In particular, relative differences with respect to AOMA Mugnaini et al. (2022) frequently exceed several tens of percent and, in some cases, are considerably larger, especially for the lower-frequency modes. This behaviour is expected, as damping estimation in operational modal analysis is well known to be more sensitive to noise, limited measurement duration, and modelling assumptions than frequency estimation. Moreover, small absolute differences in $\zeta_{n}$ can translate into large relative deviations when the reference damping level is low.
Overall, the results indicate that while the identification of $\omega_{n}$ is highly reliable and consistent across the considered approaches, the estimation of $\zeta_{n}$ remains more method-dependent. However, to fully assess the performance of NExT-LF, the identified $\bm{\phi}_{n}$ must also be examined. To this end, Table˜4 reports the diagonal MAC values between the $\bm{\phi}_{n}$ identified via NExT-LF and those obtained in Mugnaini et al. (2022).
Table 4: Diagonal terms of the MAC value matrices obtained by correlating the identified (NExT-LF, NExT-ERA, and SSi-CVA) $\bm{\phi}_{n}$ and the reference results in Mugnaini et al. (2022).
Diagonal terms of the MAC value [-] matrix
Method/Mode 1 2 3 4 5 61 7 8 9 10 111 121 13 14
NExT-LF 0.97 0.93 1.00 1.00 0.77 - 0.44 0.97 1.00 0.16 - - 0.70 0.09
NExT-ERA 1.00 - 1.00 1.00 0.99 - - 0.99 1.00 0.06 - - - -
SSI-CVA 1.00 - 1.00 1.00 0.93 - 0.99 1.00 1.00 0.97 - - 0.98 1.00
1 These $\bm{\phi}_{n}$ cannot be correlated to Mugnaini et al. (2022) as they are not reported therein.
The MAC results in Table˜4 further contextualise the $\omega_{n}$ and $\zeta_{n}$ comparisons. In particular, NExT-LF achieves very high correlation for several dominant modes (e.g., Modes 1, 3, 4, 8, and 9, with MAC values close to unity), supporting the strong agreement observed for $\omega_{n}$ and indicating that the associated $\bm{\phi}_{n}$ are consistently retrieved. At the same time, the MAC values highlight that the increased modal content identified by NExT-LF comes with a non-uniform level of shape fidelity across all modes: some modes exhibit reduced correlation with respect to Mugnaini et al. (2022), notably Mode 5 (MAC =0.77=0.77) and, more markedly, Mode 7 (MAC =0.44=0.44), as well as very low agreement for Mode 10 (MAC =0.16=0.16) and Mode 14 (MAC =0.09=0.09). This suggests that, although NExT-LF is effective in extracting a larger set of modes than NExT-ERA, a subset of the identified shapes may be more sensitive to the limited instrumentation and to the specific characteristics of the identification procedure, leading to weaker correspondence with the AOMA reference in Mugnaini et al. (2022). Conversely, SSI-CVA shows consistently high MAC values for the modes that can be compared (e.g., Mode 7 and Mode 10), indicating stronger agreement in those cases, albeit without the same increase in the number of extracted modes relative to NExT-ERA. For completeness, the 14 identified $\bm{\phi}_{n}$ are shown in Figure˜11. Note that in the cases where the MAC is more than 0.95, the NExT-LF-identified $\bm{\phi}_{n}$ are shown; otherwise, $\bm{\phi}_{n}$ from Mugnaini et al. (2022) are preferred, with an exception made for mode 12 (3rd coupled), which is retrieved from the SSI-CVA identification.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((m))
((n))
Figure 11: The 14 $\bm{\phi}_{n}$ (a-n) identified from the Airbus Helicopters H135 BMR blade AVT.
All $\bm{\phi}_{n}$ shown in Figure˜11 are consistent with the dominant motions reported in Table˜3. Nevertheless, the trajectories associated with the two pitching modes (Figures˜11(g) and 11(n)) cannot be directly inferred from the corresponding plots, as rotational degrees of freedom are not measured on the blade. These modes are therefore classified as pitching based on the FEM evidence reported in Weber et al. (2021). In addition, the three modes newly identified in the present campaign—namely, the coupled modes (Figures˜11(f), 11(k) and 11(l))—do not exhibit a clear dominant displacement direction. For this reason, they are labelled as coupled, at least in terms of combined lagging and flapping components.
5 Discussion
The results obtained for the XB-2 spar (Section˜3) and the Airbus Helicopters H135 BMR blade (Section˜4) show that NExT-LF enables accurate identification of the modal parameters $\omega_{n}$, $\zeta_{n}$, and $\bm{\phi}_{n}$. In the investigated cases, NExT-LF provides more reliable identifications than NExT-ERA, both in terms of agreement with benchmark values and robustness of the extracted modal set. This aspect is particularly relevant as both approaches rely on NExT to obtain impulse response functions from output-only data. For the helicopter blade, NExT-ERA does not retrieve the full set of modes identified by NExT-LF, whereas NExT-LF returns a richer modal content within the frequency band of interest.
Nevertheless, the helicopter blade case highlights a limitation shared by both NExT-based methods: The identification of $\bm{\phi}_{n}$ becomes less reliable above approximately $53~\mathrm{Hz}$. Beyond this frequency, the diagonal MAC values with respect to Mugnaini et al. (2022) drop below $0.7$, reaching $0.06$ (NExT-ERA) for Mode 10 and $0.09$ (NExT-LF) for Mode 14. This behaviour is likely driven by the distribution of excitation energy over the analysed bandwidth. The XB-2 spar is excited through broadband-limited random (white) noise using a shaker, whereas the H135 BMR blade identification is based on an AVT conducted in a laboratory environment, without externally applied broadband forcing (e.g., wind). The resulting excitation levels are markedly lower, which can prevent sufficient activation of the higher-frequency modes (here, above $\sim 50~\mathrm{Hz}$), ultimately reducing the spatial coherence of the estimated $\bm{\phi}_{n}$. The observed degradation therefore appears primarily NExT-dependent rather than LF-dependent. Consistently, the effect is more pronounced for NExT-ERA, which does not yield stable modes beyond $60~\mathrm{Hz}$ in this dataset. It should be mentioned that SSI-CVA does not exhibit the same limitation.
A further element is related to instrumentation and reference selection. The XB-2 spar represents a simple cantilever configuration, for which NExT can retrieve meaningful impulse response functions using a single reference channel. For the helicopter blade, two reference channels are required, as accelerations are measured along two axes. However, the absence of torsional and axial measurements may also limit the ability of NExT-based procedures to reconstruct modes with significant torsional content. In this respect, the pitching modes are not retrieved with satisfactory correlation by NExT-LF and NExT-ERA, which is consistent with the lack of measured rotational degrees of freedom.
Overall, the present results suggest that NExT-LF cab be extended from prior applications on civil structures Dessena et al. (2025) to aeronautically relevant configurations, provided that adequate attention is given to reference-channel selection and to the choice of the analysed frequency band and/or the excitation conditions adopted during testing.
6 Conclusions
In this work, the Natural Excitation Technique (NExT) is paired with the Loewner Framework (LF) to extend NExT-LF from the Operational Modal Analysis (OMA) of civil structures to that of aeronautically relevant configurations. The main outcomes of this study can be summarised as follows:
  • To the authors’ knowledge, this is the first time in the literature that NExT-LF is applied for the OMA of aeronautically relevant systems;
  • NExT-LF is applied for the OMA of the well-known eXperimental BeaRDS 2 Flexible Wing Spar Model and of an Airbus Helicopters H135 Bearingless Main Rotor Blade;
  • The natural frequencies identified by NExT-LF are coherent to those obtained with the benchmarks methods, NExT with the Eigensystem Realization Algorithm (NExT-ERA) and Stochastic Subspace Identification with Canonical Variate Analsysis (SSI), and data;
  • The identification of the damping ratios exhibits a significantly larger dispersion across methods and benchmark data, with comparatively high relative errors observed for all considered modes, particularly in the helicopter blade case;
  • The mode shapes retrieved in the lower and highly excited frequency band show a high (near 1) correlation (modal assurance criterion) with the benchmark data;
  • Three further modes are found w.r.t. the benchmark results of the helicopter blade. Two by NExT-LF and one by SSI.
[…]
Conceptualisation, G.D. and M.C.; methodology, G.D.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D.; resources, G.D., M.C. and OE.BM.; data curation, G.D..; writing—original draft preparation, G.D.; writing—review and editing, G.D., M.C. and OE.BM. visualization, G.D.; project administration, G.D.; funding acquisition, G.D., M.C. and OE.BM..
[…]
The first and third authors have been supported by the Madrid Government (Comunidad de Madrid, Spain) under the Multiannual Agreement with the Universidad Carlos III de Madrid (IA_aCTRl-CM-UC3M). The second author is supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management).
[…]
Not applicable.
[…]
Not applicable.
[…]
Data supporting this study (Modal identification of the XB-2 spar and H135 BMR blade modal parameters – Data file 1) are openly available from the Zenodo Repository at (DOI to be reserved). The flexible wing spar experimental data (Data file 2) used in this study are available in the CORD - Cranfield University research data repository entry Data supporting: ’Ground Vibration Testing of a Flexible Wing: A Benchmark and Case Study’ available at https://doi.org/10.17862/cranfield.rd.19077023. The helicopter blade experimental data (Data file 3) used in this study are not publicly available and therefore cannot be redistributed by the authors.
[…]
The authors declare no conflicts of interest.
[…]